OpenStudy (anonymous):

The probability that a random newborn is male is roughly 0.51 . What is the probability that at least 2 of 4 newborns will be male?

OpenStudy (rizags):

Bernoulli Trials?

OpenStudy (rizags):

Is this for alg 2?

OpenStudy (anonymous):

i think is Bernoulli Densities, not alg2 is probability and statistic

OpenStudy (rizags):

Ok. Either way this will work:

OpenStudy (anonymous):

ok

OpenStudy (rizags):

Suppose we have 4 trials, AT LEAST 2 of which are successes. Thus, the probability that we have 2, 3, or 4 successes is 1-P(1)=P(2,3,4). So let's calculate the probability that 1 out of 4 newborns is male. We do that like so:

OpenStudy (rizags):

\[\left(\begin{matrix}4 \\ 1\end{matrix}\right)\times0.51^1\times0.49^3=0.24\]

OpenStudy (rizags):

Finally, to get P(2, 3, or 4), We take 1-P(1):\[\huge 1-P(1)=1-0.24=0.76\]

OpenStudy (rizags):

Make sense?

OpenStudy (anonymous):

ya is does make sense, so I need to add 2 3 and 4 ?

OpenStudy (rizags):

Ok, see what we have here is a complement probability. Instead of calculating P(2)+P(3)+P(4), we just calculate P(1) and then subtract it from 1, because:\[\huge \color{blue}{P(2, 3, 4)=1-P(1)}\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so i do the samething as we didnt with p(1)? 1-p(2), 1-p(3)

OpenStudy (rizags):

no. The answer above is your final answer. There's no Further calculation

OpenStudy (anonymous):

also did u subtract .51 to get .49?

OpenStudy (rizags):

P(1)+P(2)+P(3)+P(4)=1

OpenStudy (rizags):

And yes

OpenStudy (anonymous):

so 1 2 3 4 need to add up 1 i see

OpenStudy (rizags):

Medal plz

OpenStudy (anonymous):

\[P(2) = \left(\begin{matrix}4 \\ 2\end{matrix}\right) * .59^2*.49^2\]

OpenStudy (rizags):

not 0.59 and 0.49, but 0.51 and 0.49

OpenStudy (anonymous):

mistype

OpenStudy (rizags):

Remember, for this problem, all you have to do is find P(1), then subtract from 1. THATS IT. If you want to validate then calculate P(2)+P(3)+P(4), which should give you the same answer (0.76)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got like P(2) 0.37

OpenStudy (rizags):

Yes that's correct. Now do P(3) and P(4) if you want

OpenStudy (anonymous):

p(3)=.26 and p(4)= .07

OpenStudy (rizags):

Oh no, I've made a mistake! I LEFT OUT P(0) SORRY!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so is 0 1 2 3 and 4

OpenStudy (rizags):

NEW ANSWER CORRECTED:\[\large P(0)+P(1)+P(2)+P(3)+P(4)=1\]

OpenStudy (rizags):

SO\[\large P(2)+P(3)+P(4)=1-(P(0)+P(1))\] AND\[\huge \color {red}{P(0)=0.06}\] AND \[\huge \color{blue}{P(1)=0.24}\]SO\[\large 1-(P(1)+P(2))=1-(0.06+0.24)=0.70\]

OpenStudy (rizags):

That is correct now.

OpenStudy (rizags):

Crap, mistype in response, should say 1-(P(0)+P(1))

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

somehow i see a equal number .70=.70

OpenStudy (anonymous):

found the answer is .702335

OpenStudy (rizags):

yep

OpenStudy (anonymous):

.70235